if z=(i)^(i)^(i) where i = root(-1) , then what will be the value of |z| 

Grade:11

1 Answers

Adithya
5 years ago
i= (eipi/2)i  =e-pi/2
 
So,  (i)(i)^i  =(i)e^(-pi/2)  ,Hence   |z|  =   |(i)e^(-pi/2|  =   |i|e^(-pi/2)   =  1.(Since the  |zr|  =  |z|r)
 
So Answer is 1